OpenStudy (anonymous):

The line passes through the points ( 2, 1) and ( 5, 4)

OpenStudy (imtiaz7):

do you want to find the equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is the question?

OpenStudy (solomonzelman):

Find the slope using the slope formula, \(\LARGE\color{black}{ m=\frac{y_1-y_2}{x_1-x_2} }\) Then find the y-intercept using any of your given points, and using the slope.

OpenStudy (solomonzelman):

First find the slope. Can you do this?

OpenStudy (anonymous):

Yeah!

OpenStudy (imtiaz7):

here you can use the formula to find the equation that is called: Two point form (y-y1)/(y2-y1) =(x-x1)/(x2-x1)

OpenStudy (imtiaz7):

where x1=2, y1=1 and x2=5, y2=4

OpenStudy (solomonzelman):

Close, use the point slope formula, after you know the slope. (y-y1) =m(x-x1)

OpenStudy (anonymous):

So... a. y = x + 1

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

your slope is correct y intercept has a problem

OpenStudy (anonymous):

ugh smh